OpenStudy (anonymous):

As the marketing designer for Burger King, you are given a 4 in by 6 in logo and have been told to create an advertisement with side lengths six times larger. What is the area of your advertisement (in square feet)? [12 inches = 1 ft] A. 6 ft2 B. 36 ft2 C. 72 ft2 D. 3 ft2

OpenStudy (anonymous):

first find out the lengths of the advertisement is. since it's 6 times larger it would be (4x6) inches by (6x6) inches

OpenStudy (anonymous):

Six times larger means: (4*6) and (6*6) So, the Area will be: \[Area = Length \times Breadth\] Can you go further now??

OpenStudy (anonymous):

don't forget to turn the inches into feet before finding the area

OpenStudy (anonymous):

so do i distribute?

OpenStudy (anonymous):

Length has become = (4*6) = 24 inches Breadth has become = (6*6) = 36 inches.. These both are in inches, so firstly convert them into Feet by putting 12 inch = 1 ft... Can you do it or not??

OpenStudy (anonymous):

is it 5ft ^2

OpenStudy (anonymous):

Dude you don't have to add you have to multiply the feets you are getting for length and breadth...

OpenStudy (anonymous):

Firstly tell, Length in feet = ????

OpenStudy (anonymous):

im lost i dont get what your saying

OpenStudy (anonymous):

What is the Length you are getting??

OpenStudy (anonymous):

is the answer 6

OpenStudy (anonymous):

if you turn your inches into feet then 24in becomes 2ft and 36in becomes 3ft. so the square footage would be 2^2 by 3^2 which equals 4x9

OpenStudy (anonymous):

Yes the answer is 6 feet squared..

OpenStudy (anonymous):

ok thanx for the help!!

OpenStudy (anonymous):

wait or is it 36??

OpenStudy (anonymous):

How you think it is 36??

OpenStudy (anonymous):

if you are doing square feet then it would be 36

OpenStudy (anonymous):

See, Area = Length*Breadth You have length and breadth just multiply them to get the answer..

OpenStudy (anonymous):

its 6 feet squared, sorry about that

OpenStudy (anonymous):

@triciaandchad_2009 do you know how square feet came???

OpenStudy (anonymous):

if you do that then you are not doing it in square footage you are doing it just in feet in order to do square footage then you must do the ^of the area

OpenStudy (anonymous):

Area = Length(feet)*Breadth(feet) = Length* Breadth (sq. feet) My dear friend Area is always having the units of square.. If you multiply feet with feet you will automatically get square feet...

OpenStudy (anonymous):

@Tblack have faith on yourself and go with 6..

OpenStudy (anonymous):

aha ok